# Training Bilingual LMs with Data Constraints in the Targeted Language

### v1

@@ Abstract
Large language models are trained on massive scrapes of the web, as required by current scaling laws. Most progress is made for English, given its abundance of high-quality pretraining data. For most other languages, however, such high quality pretraining data is unavailable. In this work, we study how to boost pretrained model performance in a data constrained target language by enlisting data from an auxiliary language for which high quality data is available. We study this by quantifying the performance gap between training with data in a data-rich auxiliary language compared with training in the target language, exploring the benefits of translation systems, studying the limitations of model scaling for data constrained languages, and proposing new methods for upsampling data from the auxiliary language. Our results show that stronger auxiliary datasets result in performance gains without modification to the model or training objective for close languages, and, in particular, that performance gains due to the development of more information-rich English pretraining datasets can extend to targeted language settings with limited data.
@@
Skyler Setothanks: Equal contribution, Maartje ter Hoeve11footnotemark: 1, He Bai, Natalie Schluter, David Grangier Apple {sseto,m_terhoeve,hbai22,natschluter,d_grangier}@apple.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance on many tasks, including mathematical reasoning, coding capabilities, and knowledge-based question answering (Brown et al., 2020; Bubeck et al., 2023; OpenAI, 2023) through pretraining on large corpora of web text. The vast majority of LLMs have been trained and evaluated on English, motivated by the abundance of high quality data. For other languages such large quantities of high quality and information-rich data are unavailable. Consequently, breakthroughs of similar scale are lacking for other languages, due to the lack of good data for pretraining.
Most non-English progress comes from relatively small bilingual models (e.g., Le et al., 2019; De Vries et al., 2019; Martin et al., 2019; Scheible et al., 2020; Wei et al., 2023a; Faysse et al., 2024), or larger massively multilingual models (e.g., Le Scao et al., 2023; Intrator et al., 2024; Üstün et al., 2024). Other LLMs such as Llama-2, GPT-3, and PaLM-2 that perform well across a variety of languages are trained primarily on English data with less than 20% of their data coming from other languages Xu et al. (2024). However, little progress has been made in understanding when an auxiliary language (such as English) can help learning a target language. This is particularly relevant to understand in the case of target languages for which limited data is available.
@@
This work studies the challenge of building datasets for pretraining language models with limited target language data. While much of the progress on building datasets for language model pretraining has focused on collecting and filtering more data, typically English (Figure 1(a)), we investigate whether these advances transfer implicitly to auxiliary-target language learning (Figure 1(b)), where English becomes an auxiliary language. In particular, this work investigates whether better English datasets also lead to better models in other languages. We focus on the impact of dataset size, filtering for data quality and style, and data selection for specialized information relevant to downstream evaluations (Figure 1(c)), matching recent advancements in state-of-the-art open-source English datasets such as FineWeb-EDU Penedo et al. (2024), and DataComp Li et al. (2024).
Our key findings are:
@@
    Auxiliary English data that is generated by some of the existing model-based data filtering pipelines for English can be helpful to complement limited data in a target language (Section 3.2);
@@
    When training with higher quality auxiliary data, more gains can be attributed to relevant information in the auxiliary data (Section 4.1-4.2);
@@
    Findings are not the same across multiple languages. We hypothesize that for languages that are “far” from English, gains from better English datasets do not help (Section 5);
@@
    There are limits to the size of models that are practical to pretrain with limited target data. This is because data size should scale linearly with model size Kaplan et al. (2020) and performance in the target language saturates without increasing target data regardless of increasing auxiliary data (Section 6.2).
@@ 2 Related Work
(a) Medium
(b) XL
Figure 2: Zero-shot accuracy of medium and XL models trained with higher quality English auxiliary data. Results are averaged over six eval datasets. We compare training with different auxiliary datasets (colors) on both English (solid) and German (striped) evaluations. Better English datasets show large increases in English, and smaller increases in German.
@@
While much of the LLM research has focused on English, large-scale transformer-based multilingual language models have been trained on large multilingual corpora including mBERT Pires (2019), XLM Conneau and Lample (2019), mT5 Xue (2020), PolyLM Wei et al. (2023b), and Bloom Le Scao et al. (2023). These works focus on training models that are language balanced and can reason in over 100 languages. Other state-of-the-art language models such as Llama 2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Palm 2 Anil et al. (2023) have multilingual capabilities but over 90% the training data is English, and these models are shown to perform poorly across a variety of languages, such as south east Asian languages Nguyen et al. (2023).
@@
Philippy et al. (2023) present a comprehensive survey on cross-lingual transfer in multilingual language models. The survey explains that cross-lingual transfer is a well studied topic for classic NLP tasks, such as part-of-speech tagging, named entity recognition, dependency parsing, natural language inference, machine translation, etc., although findings are not always consistent. Cross-lingual transfer is less well studied for the language modeling objective, and for modern down-stream evaluation tasks, such as ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), SCIQ Welbl et al. (2017), WinoGrande Sakaguchi et al. (2021), etc. Philippy et al. end with a number of recommendations, one of which is to study cross-lingual transfer in more detail for generative models, given their exceptional performance in recent years. Although our work does not directly study cross-lingual transfer (we do not first train on the auxiliary language, and then on the target language), our findings help understand how an auxiliary language can help a target language for which limited data is available.
@@
Selecting high quality data for pretraining LLMs remains an active area of research. Early research on data selection was based on heuristics. For example, the original GPT-2 model was pretrained on outbound links filtered from Reddit based on heuristic indicators of whether users found the links interesting, educational, or funny Radford et al. (2019). Prior approaches also upsampled documents from high quality sources like Wikipedia Gururangan et al. (2022), or used combinations of heuristics such as absence of stop words, document length, word length, etc. Rae et al. (2021). Other works select data based on quality filters and time of collection Longpre et al. (2023), model-based quality filtering Sachdeva et al. (2024); Li et al. (2024), or textbook quality knowledge Gunasekar et al. (2023); Li et al. (2023b); Kong et al. (2024). An alternative approach to data selection is re-weighting data samples to select the best data mixtures for training Fan et al. (2023); Xie et al. (2024), or importance sampling based on a downstream task Grangier et al. (2024b, a); Xie et al. (2023). Still a majority of these filtering techniques are applied to English-only datasets, and multilingual datasets such as mC4 have limited data filtering Xue (2020).
@@ 3 Using English Data Selection Pipelines to Complement Limited Target Data
Existing data selection pipelines have been shown to be effective in monolingual (English) pretraining. We investigate whether these pipelines are useful in the bilingual setup with limited target data. We always take English as the auxiliary language. Initially, we perform all experiments with German as the target language. In Section 5 we discuss how our findings extend across multiple languages.
(a) Medium
(b) XL
Figure 3: Zero-shot accuracy of medium and XL models trained with higher quality English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English and German.
@@ 3.1 General Implementation Details
We train decoder-only transformer models Vaswani et al. (2017) at different scales: medium and XL Brown et al. (2020). Models use the PolyLM tokenizer Wei et al. (2023b), with a total vocabulary size of 256K tokens using BPE. Models are trained for 30K (medium) or 100K steps (XL) with batch size 1024. Additional hyperparameters and model details are in Appendix A.
@@
We consider access to approximately 250M tokens [^1] This is chosen to simulate the amount of data that would exist in the tail of mC4. We use German for our experiments to facilitate comparisons with having additional data, and to study translation and other methods. from the target language. We upsample the 250M tokens of target language data to total 5% of the training time, and the remaining 95% are different English auxiliary datasets or mC4 German for comparisons to monolingual models. All data is pretokenized with packing to the full context length, and shuffled during training.
@@
We consider the average over six general understanding QA tasks: ARC-Easy, ARC-Challenge, Hellaswag, PIQA, SciQ, and Winogrande. These tasks are knowledge-based tasks that small models with limited data still perform well and many require knowledge. Non-English evaluations are conducted via translation of the original dataset. We use a mix of proprietary large language models to ensure good translations.
@@ 3.2 Better English Datasets
We compare the performance of models trained on combinations of German (target) and English (auxiliary) with varying existing English datasets based on the common crawl: mC4 Xue (2020), RedPajamav2 (RPJv2) Computer (2023), RefinedWeb (RFW) Penedo et al. (2023), FineWeb (FW), and FineWeb-EDU (FWE) Penedo et al. (2024). These datasets have been constructed following a pipeline of different filtering steps outlined in Figure 1(a). The resulting datasets have higher quality filtering and cover different snapshots of the common crawl. They are all aimed at general language model pretraining. Further details for the datasets are available in Appendix B. For comparison with monolingual models trained only on German (target), we consider both German mC4, and a low quality version of German mC4 referred to as “no ARC German mC4”. [^2] We downsample data based on the ARC dataset using the clustered importance sampling procedure in Section 4.2 to approximate having low quality data which is not in-domain but in the target language. This dataset serves as a baseline for a low quality dataset with less relevant data to the downstream tasks. We clarify this procedure in greater detail in the respective section as the details beyond low quality and out of domain are not critical.
@@
Across all tasks in Figure 2, we observe that for English evaluations, better quality English datasets attain substantially higher performance on the English tasks. For the medium models, this corresponds to up to 5% increase between the worst and best performing English datasets. For the XL models, the performance improvement in English is up to 9%. For the same benchmarks translated into German, the performance increase is around 1% for the medium model and 2% for the XL model. Of all compared English datasets, FineWeb-EDU achieves the best average down-stream performance, and similar performance to the no ARC German dataset. There are two primary factors we hypothesize contribute to FineWeb-EDU achieving better performance on downstream tasks: high-quality data filtering, and information filtering. Next, we investigate these factors in more detail.
(a) Medium
(b) XL
Figure 4: Zero-shot accuracy of medium and XL models trained upsampling different downstream datasets. Results are averaged over six eval datasets.
@@ 4 The Effect of Individual Data Transformations
Section 3.2 investigated the effect of using existing data filtering pipelines in the auxiliary language. Here we study individual data transformations in more detail. For all experiments, we use the same experimental setup as in Section A. Unless otherwise stated, we refer to the mC4 datasets as “Base.”
@@ 4.1 High Quality Filtering
While filtering strategies achieve strong results on English downstream evaluations, training a filtering model can require more data than available, and typical high quality datasets may not be available in the target language for filtering. Further, it is unclear to what extent better filtering strategies would improve performance in a target language that might not benefit from seeing purely higher quality data in the auxiliary language.
@@
To test the impact of filtering, we use the OH+ELI5 fast text classifier to filter data, and filter a large portion of the mC4 to the top 10% high quality documents following Li et al. (2024). Filtering with this classifier leads the model to train on high quality English, and question answer style data which specializes the model towards downstream evaluation. We compare the performance of models trained with English filtering and without, holding the German data the same for evaluation.
@@
We find in Figure 3 that for the medium models performance improvements in English evaluations are 1.5%, but for translated German (target) evaluations, there is no improvement. For the XL models, English evaluations improve by 3%, but translated German evaluations are under 1% and within 1 standard error.

### v2

@@ Abstract
Large language models are trained on massive scrapes of the web, as required by current scaling laws. Most progress is made for English, given its abundance of high-quality pretraining data. For most other languages, however, such high quality pretraining data is unavailable. In this work, we study how to boost pretrained model performance in a target language with insufficient pretraining data for training a high performing language model, by enlisting data from an auxiliary language for which high quality data is available. We study this by quantifying the performance gap between training with data in a data-rich auxiliary language compared with training in the target language, exploring the benefits of translation systems, studying the limitations of model scaling when data is limited in the target languages, and proposing new methods for upsampling data from the auxiliary language. Our results show that stronger auxiliary datasets result in performance gains without modification to the model or training objective for close languages, and, in particular, that performance gains due to the development of more information-rich English pretraining datasets can extend to targeted language settings with limited data.
@@
Skyler Setothanks: Equal contribution, Maartje ter Hoeve11footnotemark: 1, Richard He Bai, Natalie Schluter, David Grangier Apple {sseto,m_terhoeve,richardbai,natschluter,d_grangier}@apple.com
@@ 1 Introduction
The abundance of high quality English data has given rise to exceptional language modeling performance in English (Brown et al., 2020; Bubeck et al., 2023; OpenAI, 2023). However, there are many other languages for which a reasonably large amount of data is available, yet not in equally large amounts as for English. Taking mC4 Xue (2020) as an example, languages like German, French, Chinese, and Japanese have $\sim$10-100 times less data than English. Consequently, most non-English progress comes from relatively small bilingual models (e.g., Le et al., 2019; De Vries et al., 2019; Martin et al., 2019; Scheible et al., 2020; Wei et al., 2023a; Faysse et al., 2024), or larger massively multilingual models (e.g., Le Scao et al., 2023; Intrator et al., 2024; Üstün et al., 2024). Other LLMs such as Llama-2, GPT-3, and PaLM-2 that perform well across a variety of languages are trained primarily on English, with less than 20% of data from other languages Xu et al. (2024).
Several works study cross-lingual transfer, but are limited to finetuning smaller models ($\sim$100M parameters) with limited data, or classical NLP tasks, such as part-of-speech tagging, named entity recognition, and natural language inference (e.g., Chang et al., 2023; de Vries et al., 2022; Faisal and Anastasopoulos, 2024). They do not evaluate pretrained language models on contemporary knowledge-based downstream tasks. It is still unclear how to optimally make use of bi- or multilingual data, to optimally increase performance in a target language. This is especially relevant in data constrained settings. For example, imagine one has access to only $500$K documents in a target language, and wants to pretrain a language model for this language. $500$K documents is sufficient to attain a reasonable perplexity score, but this amount of data is insufficient to do well on modern knowledge-based downstream tasks. Many languages in the tail of mC4 fall in this category. Currently, a systematic analysis of how to use an auxiliary language in this scenario is lacking, leaving practitioners to a trial-and-error based approach, mostly based on intuition.
In this work, we provide such a systematic analysis. In particular, we are interested in whether “higher quality” auxiliary data also leads to higher model performance in the target language. We choose English as the main auxiliary language, matching a common practical setting. To investigate the effect of linguistic similarity, we also experiment with Chinese as the auxiliary language, although being somewhat bottlenecked by the amount of available data in Chinese.
@@
To measure the impact of “better” auxiliary data, we explore well-tested filtering techniques for a monolingual, English, setting. We focus on the impact of dataset size, filtering for data quality and style, and data selection for specialized information relevant to downstream evaluations (Figure 1(c)), matching recent advancements in state-of-the-art open-source English datasets such as FineWebEDU Penedo et al. (2024), and DataComp Li et al. (2024) (Figure 1). Currently, it is still unclear how these techniques extend to a bilingual setting; as we confirm in our experiments, not every filtering technique that performs well in a monolingual setting performs well in a bilingual setting. Specifically, our key findings are:
@@
    Auxiliary English data that is generated by some of the existing model-based data filtering pipelines for English can be helpful to complement limited data in a target language (§3.2);
@@
    Findings are not the same across multiple target languages. We hypothesize that for languages that are “far” from English, gains from better English datasets do not help (§3.3);
@@
    When training with higher quality auxiliary data, slightly more gains can be attributed to relevant information in the auxiliary data (§4.1-4.2);
@@
    There are limits to the model sizes that are practical to pretrain with limited target data: data size should scale linearly with model size Kaplan et al. (2020), and performance in the target language saturates without increasing target data, regardless of increasing auxiliary data (§5.2).
@@ 2 Related Work
While much of the LLM research has focused on English, large-scale transformer-based multilingual language models have been trained on large multilingual corpora including mBERT Pires (2019), XLM Conneau and Lample (2019), mT5 Xue (2020), PolyLM Wei et al. (2023b), and Bloom Le Scao et al. (2023). These works focus on training models that are language balanced and can reason in over 100 languages. Other SOTA language models such as Llama 2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Palm 2 Anil et al. (2023) have multilingual capabilities, but over 90% the training data is English, and these models perform poorly across a variety of languages, such as south east Asian languages Nguyen et al. (2023).
@@
Philippy et al. (2023) present a comprehensive survey on cross-lingual transfer in multilingual language models. The survey explains that cross-lingual transfer is a well studied topic for classic NLP tasks, such as part-of-speech tagging, named entity recognition, dependency parsing, machine translation, etc., although findings are not always consistent. Cross-lingual transfer is less well studied for the language modeling objective, and for modern downstream evaluation tasks, such as ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), SCIQ Welbl et al. (2017), WinoGrande Sakaguchi et al. (2021), etc. Philippy et al. end with a number of recommendations, one of which is to study cross-lingual transfer in more detail for generative models, given their recent exceptional performance. Our work differs from prior work on cross-lingual transfer in three ways: (i) we do not directly study cross-lingual transfer as we do not first train on the auxiliary language, and then on the target language, (ii) we pretrain 1B and 3B parameters models that are 10-30x the size of typical BERT models, and (iii) instead of finetuning, we focus on how models learn to share between languages during pretraining.
@@
Selecting high quality data for pretraining LLMs remains an active area of research. Early research on data selection was based on heuristics. For example, the original GPT-2 model was pretrained on outbound links filtered from Reddit, based on heuristic indicators of whether users found the links interesting, educational, or funny Radford et al. (2019). Prior approaches also upsampled documents from high quality sources like Wikipedia Gururangan et al. (2022), or used combinations of heuristics such as absence of stop words, document length, word length, etc. Rae et al. (2021). Other works select data based on quality filters and time of collection Longpre et al. (2023), model-based quality filtering Sachdeva et al. (2024); Li et al. (2024), or textbook quality knowledge Gunasekar et al. (2023); Li et al. (2023b); Kong et al. (2024). An alternative approach to data selection is re-weighting data samples to select the best data mixtures for training Fan et al. (2023); Xie et al. (2024), or importance sampling based on a downstream task Grangier et al. (2024b, a); Xie et al. (2023). Still a majority of these filtering techniques are applied to English-only datasets, and multilingual datasets such as mC4 have limited data filtering Xue (2020).
@@ 3 Using English Data Selection Pipelines to Complement Limited Target Data
Existing data selection pipelines have been shown to be effective in monolingual (English) pretraining. We investigate whether these pipelines are useful in the bilingual setup with limited target data. In §3.2, we report our findings for experiments with English as the auxiliary language and German as the example target language. In §3.3 we discuss how our findings extend across multiple languages.
@@ 3.1 General Implementation Details
We train decoder-only transformer models Vaswani et al. (2017) with 1.3B parameters. Models use the PolyLM tokenizer Wei et al. (2023b), with a total vocabulary size of 256K tokens using BPE. Models are trained for 100K steps with batch size 1024. Additional hyperparameters and model details are in Appendix A. We report results for the 1.3B model in the main body, and refer to Appendix G for results on a 300M model trained for 30K steps. These model sizes are chosen as they provide reasonable (above random) performance on several benchmark QA tasks, and are commonly used for benchmarking and ablating pretraining of language models Penedo et al. (2023, 2024).
@@
To meaningfully and systematically study a data constrained scenario, we impose the following trade-off: (i) the number of tokens in the target language needs to be sufficient to start training a language model in the target language (i.e., perplexity needs to go down sufficiently), and (ii) the number of tokens in the target language should be small enough to expect meaningful differences from the auxiliary language. Through empirical exploration, we choose to include 250M tokens in the target language. Not only does this choice satisfy the trade-off posed above, but it also simulates a realistic scenario, as 250M tokens is representative for the amount of data that exists in the tail of mC4 ($\sim$20 languages). For our experiments, we choose languages with more than 250M tokens available in mC4. This allows us to compare with a monolingual setting, to better assess the impact of the auxiliary language.
We consider three baselines: (i) Target (S): 250M tokens from the target language, (ii) Target (L): enough target language for 1 training epoch, and (iii) Auxiliary (L): enough auxiliary language for 1 training epoch. For the other experiments, we consider $\sim$250M tokens from the target language, as stated above, which is repeated to account for 5% of the training steps. The remaining 95% of the training steps consists of the auxiliary data. We refer to the small 250M target language data as “(S)” and to the large (auxiliary) data pool as “(L)”.
@@
We consider the average over six general understanding QA tasks: ARC-Easy, ARC-Challenge, Hellaswag, PIQA, SciQ, and Winogrande. These are knowledge-based tasks that small models with limited data still perform well on. Non-English evaluations are conducted via translation of the original dataset for which we use a mix of proprietary large language models. Additional details are provided in §B.2.
@@ 3.2 Better English Datasets
We first compare the performance of models trained on combinations of German (target) and English (auxiliary) with varying existing English datasets based on the common crawl: mC4 Xue (2020), RedPajamav2 (RPJv2) Computer (2023), RefinedWeb (RFW) Penedo et al. (2023), and FineWebEDU (FWE) Penedo et al. (2024). These datasets have been constructed following a pipeline of different filtering steps outlined in Figure 1(a). The resulting datasets are of higher quality and cover different snapshots of the common crawl. Further details are available in Appendix B.
@@
We report results in Figure 2. For English evaluations, better quality English datasets attain substantially higher performance on the English tasks (up to 9%). For the same benchmarks translated into German, the performance increase is around 2%. Of all compared English datasets, FineWebEDU achieves the best average down-stream performance, and within 1% of the DE (L) comparison. There are two primary factors we hypothesize contribute to FineWebEDU achieving better performance on downstream tasks: high-quality data filtering, and relevant information filtering. We investigate this in more detail in §4.
3.3 Experiments Across Multiple Languages
Figure 2: Zero-shot accuracy of models trained with higher quality English auxiliary data. Results are averaged over six eval datasets. We compare training with different auxiliary datasets on English and German evaluations. Better English datasets show large increases in English and smaller increases in German.
Figure 3: Average zero-shot accuracy in the target language summarized for eight languages. Models trained on 100B tokens. Comparisons between a small and large amount of monolingual data from the target language, a small amount of data from the target language and mC4 English data (same distribution), and a small amount of data from the target language and FineWebEDU.
Motivation.
We add seven languages, across four language families: French, Italian, Portuguese, Spanish (Indo-European, same as German), Chinese (Sino-Tibetan), Japanese (Japonic), and Korean (Koreanic) Lewis et al. (2015). We choose these languages because of their variety in language families, the amount of data available in mC4, and the access to translated evaluation data.
Methodology.
We train with approximately 250M tokens [^1] Note that data distributions from the $\sim$250M data in mC4 looks similar for each language and are not specialized to any particular domain as evidenced by the cluster distribution visualized in Appendix E. from the mC4 corpus in the respective language and denote these models as “target language (S)”. The monolingual models are denoted as “target language (L)”. The Chinese and Korean mC4 corpora contain fewer than 100B tokens and thus the data is repeated for multiple epochs in the base language experiments. Based on our previous results, we use FineWebEDU as the auxiliary language source. We also conduct additional experiments with Chinese as the auxiliary language and Japanese as the target language, to investigate the effect of linguistic similarity. Although beyond the immediate scope of our investigation, we also investigate a multilingual setup (instead of bilingual) in Appendix H.3.
Findings.
We report results for multiple target languages and Chinese auxiliary data in Figure 3-4. We only observe improvements from FineWebEDU for the Indo-European languages in our set, with English as the auxiliary language (Figure 3). Figure 4 shows better performance for Chinese instead of English FineWebEDU as the auxiliary language for Japanese. This is a first indication of the applicability of using different auxiliary languages, based on linguistic similarity.
Figure 4: Average accuracy over zero-shot benchmark tasks in translated Japanese, comparing Chinese and English auxiliary data.
@@ 4 The Effect of Individual Data Transformations
In §3.2 we investigated the effect of using existing data filtering pipelines in the auxiliary language, but this does not isolate which components of the pipeline provide the best downstream improvements. Here, we study individual data transformations that comprise the training pipeline, including model based quality filtering and educational content upsampling, as well as transformations on top of auxiliary data, such as translation. For all experiments, we use the same experimental setup and baselines as in §3, and we focus on German as the target and English as the auxiliary language.
@@ 4.1 High Quality Filtering
Prior work shows that model-based filters and higher quality data lead to better performance in English. A key component of FineWeb is custom filters based on text quality Penedo et al. (2024), and DCLM models are trained on data which has been highly filtered based on heuristic and model-based filters Li et al. (2024). To isolate the impact high quality auxiliary data can have on target language performance, we apply a recent model-based quality filter on top of mC4 English. We choose this filter as it aims primarily at stylistic quality, and filtering with this model leads to better performance than training with RFW and FWE Li et al. (2024), two of the best performing datasets in Figure 2. Note that while filtering strategies achieve strong results on English downstream evaluations, training a filtering model can require more data than available, and high quality datasets may not be available in the target language.
@@
To test the impact of quality filtering, we use the DCLM fast text classifier. This filters data that is aimed at instruction following and high scoring posts in r/ExplainLikeImFive, and was found to be the best quality filter over RFW and other model-based filters Li et al. (2024). We compare the performance of models trained with English filtering and without, holding the German data the same for evaluation. We refer to this as DCLM Filter (following Li et al. (2024)) and filter to the top 10% of mC4.
@@
The results are in Figure 5. English evaluations improve by 3%. Translated German evaluations are under 1%, within 1 standard error.
Figure 5: Zero-shot accuracy of models trained with model based filtering of English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English and German.